OpenStudy (anonymous):

Rick was given a prescription of antibiotics that were 500 milligram tablets. He was instructed to take 1 tablet a day. Suppose he had been taken the medicine for awhile (same time every day). The amount of medicine in his body (in milligrams) after n doses: 500+500(.3)+500(.3)^2+500(.3)^3+... +500(.3)^(n-1) Find the amount of medicine that will be in his system if he follows doctor's orders. Round answer to nearest whole milligram.

OpenStudy (anonymous):

714.28 ~ 714.3

OpenStudy (anonymous):

I don't quite understand?

OpenStudy (anonymous):

it basically forms a geometric progression 500 = a r = .3 the summation of terms n of g.p is 1 - r^n sum = a x ------ 1 - r

OpenStudy (anonymous):

now for the 1 - r^n r = .3 n=500 r^n is very small so 1 - .3^500 ~ 1 so the formula for this question becomes sum = a/ ( 1 -r ) r = .3 1 - r = .7 sum = 500/.7 = 5000/7 ~ 714.3

OpenStudy (anonymous):

do you understand now?

OpenStudy (anonymous):

What does ~ stand for?

OpenStudy (anonymous):

~ is for nearly equal to

OpenStudy (anonymous):

like 7.47 = 7.47 ~ 7.5

OpenStudy (anonymous):

So the amount of medicine would be ~ 714.3 milligrams?

OpenStudy (anonymous):

yeah : )

OpenStudy (anonymous):

Thank you so much ishaan! :)